# TAPIP3D: Tracking Any Point in Persistent 3D Geometry

### v2

@@ 4.2 Ablations
In Table 5(a), TAPIP3D that tracks in XYZ (world) and XYZ (Cam) space outperform versions that track in UV+D and UV+log(D) space, the latter two commonly adopted in existing state-of-the-art methods [18, 28]. Specifically, tracking in world XYZ coordinates performs best (AJ metric). Tracking in world coordinates generalizes effectively to scenarios with significant camera motion, such as those in LSFOdyssey [37].
Considering support points during attentions helps. In Table 5(b), we compare the proposed Neighborhood-to-Neighborhood Attention mechanism against conventional point-to-neighborhood cross attention. Specifically, we replace the Neighborhood-to-Neighborhood Attention, which groups each query point with its neighboring context points in both spatial dimensions, with standard 2D attention that considers only the individual query points. Our experiments show that Neighborhood-to-Neighborhood Attention significantly improves the APD metric from 75.1 to 84.1. This region-to-region cross-attention approach notably mitigates matching ambiguities by incorporating richer context compared to the simpler point-to-neighborhood matching baseline. Similar conclusion was reached in LocoTrack [5] but for 2D image plane using CNNs, which we extend to 3D feature clouds using cross-attention.
Limitations
Despite achieving state-of-the-art results, TAPIP3D’s performance declines when provided depth inputs are low-resolution or of poor quality, indicating a sensitivity to input depth fidelity. This limitation can be mitigated via depth map pre-processing methods such as completion and noise filtering. With ongoing advancements in 3D vision reconstruction models offering increasingly accurate priors, we anticipate continued improvements in the robustness of our method.
@@ 5 Conclusion
In this supplementary material, Section A provides further training details of our model, including a memory design that significantly helps reduce VRAM usage. Section B explains the details of our evaluation setup. In Section C, we present additional visualization of our method using stereo depth, enabled by FoundationStereo [43], a powerful stereo matching method. Section D provides further ablation results on the number of neighbors to offer insights into model behavior. Additionally, in Section E, we report the inference speed of compared to existing 3D tracking approaches. For extensive visual comparisons of our results in Section F, please refer to our project page.
@@ Appendix A Additional Training Details
Table 6 lists the hyperparameters used during training. We adopt all the 2D data augmentations from CoTracker3 [17], along with a smoothly varying sequence of random rigid transformations applied on the input point clouds to improve camera-centric performance. During training, we perform iterative refinements per window. The loss at the -th iteration is weighted by , where is the discount factor. The total loss for each sample is computed as the sum of these discounted losses across all iterations and windows. To mitigate unstable gradients in the early training stage, we additionally scale the total loss by 0.005.
@@ Appendix B Evaluation Details
For evaluation, we use the metrics defined in TAPVid [7] and TAPVid-3D [19] throughout the paper. All baselines, as well as our method, operate in a window-by-window manner, so they can only predict trajectories after the query frame (i.e., in the forward direction). However, the TAPVid-3D metric also requires evaluating trajectories before the query frame (i.e., in the backward direction). Therefore, to obtain these backward trajectories, we reverse the video in time and run inference again.
Following previous 2D tracking methods [17, 8], we augment the provided query points with additional support queries sampled from a 2D grid in the first frame. These points are lifted into 3D and inferred jointly with the original queries. The grid resolution is set to for all models. For datasets with known camera intrinsics, we use the ground-truth intrinsics both to lift image features into 3D feature clouds and to convert the outputs of UVD-space models into XYZ coordinates. We do not apply the trick of treating low-confidence point as occluded, as done in CoTracker3 [17].
@@
Appendix C Inference with Stereo Videos
While our method requires reasonably accurate unprojected point clouds as input, obtaining high-quality depth maps remains challenging in many real-world scenarios. To mitigate the limitation, we demonstrate how our method can be combined with FoundationStereo [43], a recent advancement in stereo matching, to produce metric-scaled point tracks from in-the-wild stereo videos.
Specifically, we follow the pipeline proposed in Stereo4D [14] to download stereo videos in the VR180 format from YouTube and rectify them. We then estimate disparity maps frame by frame using FoundationStereo [43], using them to unproject the image features into camera-centric spatio-temporal feature clouds. Figure 5 shows two examples produced by this process. Comparison videos are provided in the supplementary material (TAPIP3D-supp.mp4). While the examples are camera-centric tracking results, one can also combine the estimated disparity with camera poses from methods such as MegaSaM [23] to track in world-centric coordinates.
Appendix D Ablation on Neighborhood Size
We train our model with , where denotes the number of neighbors used in the N2N attention module, and evaluate its inference performance under different values on both DexYCB and LSFOdyssey. As shown in Table 7, the performance of TAPIP3D improves as increases, though the gain saturates when increases from 32 to 48 on LSFOdyssey. The computational complexity of the N2N attention module, given by , makes it a major bottleneck in our model, where is the number of frames, is the number of tracked points, and is the feature dimension. We set as a trade-off between performance and efficiency.
@@ Appendix D Ablation on Neighborhood Size
Appendix E Inference Speed Comparison
Table 8 compares the inference speed of our model with other feed-forward 3D tracking approaches. All speed evaluations are conducted using BF16 mixed precision on a single L40S GPU. We evaluate on 32-frame videos, with 1024 query points given in the first frame. While our method is approximately 1.3 slower than SpaTracker [44], it remains significantly faster than optimization-based approaches such as Shape-of-Motion [39] which needs hours to train on a sequence of 300 frames.

### v3

@@ 4.2 Ablations
In Table 5(a), TAPIP3D that tracks in XYZ (world) and XYZ (Cam) space outperform versions that track in UV+D and UV+log(D) space, the latter two commonly adopted in existing state-of-the-art methods [cotracker, delta]. Specifically, tracking in world XYZ coordinates performs best (AJ metric). Tracking in world coordinates generalizes effectively to scenarios with significant camera motion, such as those in LSFOdyssey [scenetracker].
Considering support points during attentions helps. In Table 5(b), we compare the proposed Neighborhood-to-Neighborhood Attention mechanism against conventional point-to-neighborhood cross attention. Specifically, we replace the Neighborhood-to-Neighborhood Attention, which groups each query point with its neighboring context points in both spatial dimensions, with standard 2D attention that considers only the individual query points. Our experiments show that Neighborhood-to-Neighborhood Attention significantly improves the APD metric from 75.1 to 84.1. This region-to-region cross-attention approach notably mitigates matching ambiguities by incorporating richer context compared to the simpler point-to-neighborhood matching baseline. Similar conclusion was reached in LocoTrack [locotrack] but for 2D image plane using CNNs, which we extend to 3D feature clouds using cross-attention.
5 Limitations
Despite achieving state-of-the-art results, TAPIP3D’s performance can be affected by the fidelity of the input depth maps. As a model that works in XYZ space, it expects the input geometry to remain stable and consistent across frames to ensure well-defined 3D trajectories. While MegaSaM generally provides reliable geometry, failures may occur in scenes with extreme depth variations or small, distant elements that appear blurred. In such cases, depth flickering or incorrect surface connections may arise, leading to degraded tracking quality. Furthermore, when high-quality depth (e.g., from sensors) is not available, TAPIP3D may perform worse on 2D metrics than UVD-space or purely 2D tracking methods, since its 2D trajectories are projected from 3D estimates and rely on geometrically consistent data. This limitation can be mitigated via depth map pre-processing methods such as completion and noise filtering. With ongoing advancements in 3D vision reconstruction models offering increasingly accurate priors, we anticipate continued improvements in the robustness of our method.
@@ 6 Conclusion
Acknowledgements
This material is based on work supported by an NSF Career award, an Amazon faculty award, and AFOSR Grant FA9550-23-1-0257.
@@ Acknowledgements
In this supplementary material, Section A provides further training details of our model, including a memory design that significantly helps reduce VRAM usage. Section B explains the details of our evaluation setup. In Section C, we provide details regarding the inference speed of our model. In Section D, we present additional visualization of our method using stereo depth, enabled by FoundationStereo [wen2025foundationstereozeroshotstereomatching], a powerful stereo matching method. Section E provides further ablation results to offer insights into model behavior. Additionally, we report the evaluation results of our model on TAPVid-DAVIS [tapvid] in Section G. For extensive visual comparisons of our results in Section H, please refer to our project page.
@@ Appendix A Additional Training Details
Table 6 lists the hyperparameters used during training. We adopt all the data augmentations from PIPs [harley2022particle] (also used in CoTracker3 [cotracker3]), including random Gaussian blur, random occlusion, random cropping, and color jittering, along with a smoothly varying sequence of random rigid transformations applied on the input point clouds to improve camera-centric performance. During training, we perform iterative refinements per window. The loss at the -th iteration is weighted by , where is the discount factor. The total loss for each sample is computed as the sum of these discounted losses across all iterations and windows. To mitigate unstable gradients in the early training stage, we additionally scale the total loss by 0.005.
@@ Appendix B Evaluation Details
For evaluation, we use the metrics defined in TAPVid [tapvid] and TAPVid-3D [tapvid3d] throughout the paper. All baselines, as well as our method, operate in a window-by-window manner, so they can only predict trajectories after the query frame (i.e., in the forward direction). However, the TAPVid-3D metric also requires evaluating trajectories before the query frame (i.e., in the backward direction). Therefore, to obtain these backward trajectories, we reverse the video in time and run inference again.
Following previous 2D tracking methods [cotracker3, tapir], we augment the provided query points with additional support queries sampled from a 2D grid in the first frame. These points are lifted into 3D and inferred jointly with the original queries. The grid resolution is set to for all models. For datasets with known camera intrinsics, we use the ground-truth intrinsics both to lift image features into 3D feature clouds and to convert the outputs of UVD-space models into XYZ coordinates. We do not apply the trick of treating low-confidence point as occluded, as done in CoTracker3 [cotracker3].
@@
Appendix C Inference Speed
To assess the efficiency of our model, we measure its inference speed on a single L40S GPU using BF16 mixed precision. We evaluate the model on 32-frame video sequences, where 1,024 query points are provided in the first frame. Our model achieves an inference speed of 11.3 FPS. It is approximately 1.3 slower than SpaTracker [SpatialTracker] on the same videos. It is of course substantially faster than optimization-based methods such as Shape-of-Motion [wang2024shape], which require several hours of inference-time optimization for a 300-frame sequence.
For monocular videos, we find that generating depth maps and estimating camera poses are the most time-consuming components in our pipeline. For instance, processing a sequence from DexYCB takes 121.1s for geometry estimation with MegaSaM, yet our model only requires 6.4s for point tracking. Among this, feature extraction accounts for 0.11s, encoding trajectory tokens for 6.04s, and the updating transformer for 0.25s. We also observe that our model is more robust to degraded camera pose quality than to degraded depth estimation. Consequently, one potential way to accelerate the pipeline is to use a more efficient depth estimator, such as CUT3R [cut3r], especially in scenarios where sensor depth is available. We expect that future improvements in depth estimation models will help reduce this bottleneck.
Appendix D Inference with Stereo Videos
While our method requires reasonably accurate unprojected point clouds as input, obtaining high-quality depth maps remains challenging in many real-world scenarios. To mitigate the limitation, we demonstrate how our method can be combined with FoundationStereo [wen2025foundationstereozeroshotstereomatching], a recent advancement in stereo matching, to produce metric-scaled point tracks from in-the-wild stereo videos.
Specifically, we follow the pipeline proposed in Stereo4D [jin2024stereo4d] to download stereo videos in the VR180 format from YouTube and rectify them. We then estimate disparity maps frame by frame using FoundationStereo [wen2025foundationstereozeroshotstereomatching], using them to unproject the image features into camera-centric spatio-temporal feature clouds. Figure 5 shows two examples produced by this process. Comparison videos are provided in the supplementary material (TAPIP3D-supp.mp4). While the examples are camera-centric tracking results, one can also combine the estimated disparity with camera poses from methods such as MegaSaM [megasam] to track in world-centric coordinates.
Appendix E Additional Ablations and Analyses
E.1 Ablation on Neighborhood Size
We train our model with , where denotes the number of neighbors used in the N2N attention module, and evaluate its inference performance under different values on both LSFOdyssey and Dynamic Replica. As shown in Table 7, the performance of TAPIP3D improves as increases, though the gain saturates when increases from 32 to 48 on LSFOdyssey. The computational complexity of the N2N attention module, given by , makes it a major bottleneck in our model, where is the number of frames, is the number of tracked points, and is the feature dimension. We set as a trade-off between performance and efficiency.
@@ E.1 Ablation on Neighborhood Size
E.2 Robustness to Image Blur
Real-world images of dynamic objects often exhibit blurriness, especially under low resolution and fast motion. To assess our model’s robustness to such degradation, we evaluate TAPIP3D‑world + MegaSaM on the LSFOdyssey dataset using synthetic Gaussian blur of varying strength. For each frame, the input image is convolved with a Gaussian kernel of standard deviation . Since image blur affects the reliability of depth estimation, we re‑estimate the depth maps of blurred images using MegaSaM.
| (px) | TAPIP3D (ours) | DELTA | ||
| APD | APD | APD | APD | |
| 0.0 | 29.2 | 87.5 | 28.0 | 90.2 |
| 0.5 | 21.2 | 87.4 | 20.3 | 90.3 |
| 1.0 | 20.6 | 86.6 | 19.8 | 90.3 |
| 2.0 | 18.1 | 86.0 | 17.5 | 90.2 |
As shown in Table 8, our model consistently achieves higher 3D accuracy than the baseline DELTA across all blur levels. We observe that even under heavy blur, the predicted 3D points remain well aligned with scene geometry, although the depth estimated by MegaSaM becomes less consistent across frames.
E.3 Robustness to Camera Pose Errors
To evaluate our model’s robustness against inaccurate camera poses, we synthetically perturb the ground-truth camera orientations. For each frame, we apply a rotation around a random axis, where the rotation angle is sampled from a zero-mean Gaussian distribution with a standard deviation of .
As reported in Table 9, our model, TAPIP3D-world, demonstrates considerable resilience to these perturbations. The performance degrades gracefully as the noise level increases, and our model consistently outperforms the DELTA baseline even under the perturbation of .
| (deg) | Method | AJ | APD | OA |
| N/A | DELTA | 37.1 | 50.1 | 82.3 |
| N/A | TAPIP3D-camera | 68.3 | 83.2 | 86.2 |
| 0.0 | TAPIP3D-world | 72.2 | 85.8 | 86.9 |
| 0.5 | TAPIP3D-world | 70.4 | 84.6 | 86.8 |
| 1.0 | TAPIP3D-world | 67.8 | 82.7 | 86.4 |
| 2.0 | TAPIP3D-world | 64.9 | 80.3 | 85.7 |
Appendix F Evaluation with Sparse Depth
In some real-world scenarios, we only have access to sparse depth measurements (e.g., from LiDAR sensors) instead of dense depth maps. To assess the usefulness of our approach under such conditions, we construct a benchmark consisting of 50 sequences from the Waymo Open Dataset [waymo] (a subset of TAPVid-3D), paired with depth map annotations obtained by densifying sparse LiDAR measurements using a simple nearest-neighbor interpolation, following the approach in DriveTrack [balasingam2024drivetrack].
We report the performance of TAPIP3D-camera, DELTA, and CoTracker3 on this benchmark in Table 10. TAPIP3D-camera achieves the best 3D accuracy among the compared methods. These results demonstrate that our method is highly usable and competitive when only sparse depths are available.
| Method | AJ | APD | AJ | APD |
| CoTracker3 | 16.5 | 24.1 | 63.7 | 78.8 |
| DELTA | 21.0 | 30.8 | 71.4 | 85.0 |
| TAPIP3D-camera | 21.8 | 31.6 | 67.7 | 82.2 |
Appendix G TAP-Vid Evaluation
For reference, we also evaluated TAPIP3D‑world on TAP‑Vid‑DAVIS [tapvid], a 2D point tracking benchmark. TAPIP3D‑world achieves an AJ of 58.9, an APD of 71.2, and an OA of 89.6, which are lower than those of recent state‑of‑the‑art 2D trackers.
We note that TAPIP3D is not designed to achieve state‑of‑the‑art 2D tracking performance, and the 2D tracks obtained by projecting 3D trajectories to 2D cannot reach sub‑pixel accuracy in monocular settings. The performance is further affected by MegaSaM’s failures on some of sequences, which lead to severely noisy depth estimates and unstable point clouds.
Appendix H 3D Track Visualization
For extensive 3D visualizations of our results, please refer to the accompanying project page, where we compare our method with the baselines SpaTracker [SpatialTracker], CoTracker3 [cotracker3] and DELTA [delta].